OpenStudy (anonymous):

ROn invested some money at 5% and $500 more at 7%. Hos total annual interest income was $215. How much did Ron invest at 5%?

OpenStudy (anonymous):

did he spend 500 at 7% or 500 MORE than what he spent at 5%? If he only spent 500 at 7%: .05x + .07(500) = 215 .05x + 35 = 215 .05x = 180 x = 3600 If he spent 500 MORE than that of 5%: .05x + .07(x + 500) = 215 .05x + .07x + 35 = 215 .12x = 180 x = 1500